OpenStudy (sdlrxo):

A jumbo chocolate bar with a rectangular shape measures 14 centimeters in​ length, 9 centimeters in​ width, and 2 centimeters in thickness. Due to escalating costs of​ cocoa, the management has decided to reduce the volume of the bar by 10​%. To accomplish this​ reduction, management decides that the new bar should have the same 2 centimeter​ thickness, but the length and width of each should be reduced by an equal number of centimeters. What should be the dimensions of the new candy​ bar?

OpenStudy (herpderp):

14 x 9 x 2 = 252 < Original Volume 252 * .9 = New volume New Volume is 226.8 (14-x) * (9-x) * 2 = 226.8 2(126 - 14x -9x + x^2) = 226.8 2x^2 -28x - 18x +252 = 226.8 2x^2 - 46x + 25.2 = 0 Use the quadratic formula to solve. (Deleted my answer on accident)